OpenStudy (anonymous):

Pre Calc: Which statement best describes a method that can be used to sketch the graph y = |x – 2|? A.)Translate the graph of y = |x| two units left. B.)Translate the graph of y = |x| two units down. C.)Translate the graph of y = |x| two units right. D.)Translate the graph of y = |x| two units up. Thanks so much <3 xoxo

OpenStudy (calculusfunctions):

If given y = f(x - d), then y = f(x) is translated d units left if d < 0 which gives (x + d), or y = f(x) is translated d units right if d > 0 which gives (x - d)

OpenStudy (anonymous):

@calculusfunctions Is the answer B?